# Learned Contextual LiDAR Informed Visual Search in Unseen Environments

### v5

@@ III-E2 Viewpoint Selection
Candidate viewpoints are selected based on a computationally efficient heuristic utility function. High value viewpoints are those that are 1) near the robot, 2) allow inspection of many unknown map points and 3) result in inspection of non-map points. This process is shown in Fig. 1(f). The utility functions is a weighted sum of the following:
  • Penalize distance from robot position to viewpoint
  • Reward expected number of unknown cells discovered
  • Reward viewpoints with frontiers near the path to the candidate viewpoint
  • Reward viewpoints that result in inspection of a scaled number non-map points
Scaling parameters on each term are tuned during experiments. Weight are set such that the agent is inclined to prioritize non-map points and those viewpoints that result in high number of expected cells discovered. The next waypoint is selected by sending the highest utility viewpoint, $x^{*}_{\text{next}}$, to the robot.
@@ IV Results
We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed successfully on real robot hardware in unseen indoor environments.
IV-A LiDAR Scan Classification
First we compare the learned classification policy to the ground truth classifier. The output of the policy is a vector of size $n_{i}$ normalized between -1 and 1 for each scan point. If this value is non-negative, the point is classified as a member of $\mathcal{E}^{\text{map}}$. Otherwise the point is classified as a member of $\mathcal{E}^{\text{non-map}}$. We express performance in terms of per-scan accuracy of the classified LiDAR scans from the neural network estimator versus ground truth scan classification from map information. This metric for a classified scan $s^{\text{class}}_{i}$ at time $i$ compared to ground truth $s^{\text{class, true}}_{i}$ is given by
IV-B Simulation and Baseline Comparison
Simulations are performed using Gazebo-ROS on the Turtlebot [32]. ROS move base is used to navigate to the goal points given from the viewpoint selection methods compared [25]. Baselines are: 1) a lightly modified version of [42] (Multisensor Frontier Exploration), 2) the Next-Best View Planner (NBVP) [5], 3) a Rapidly-Exploring Random Tree (RRT) [41], and 4) LIVES given ground-truth scan classifications. Ref. [42] described a highly efficient planner for multisensor 3D exploration in underwater environments that has also been extended to multi-drone exploration [3]. The modifications from the original paper are as follows. First, the method is projected into 2D. Second, candidate viewpoint selection is changed. Originally, candidate viewpoints are selected surrounding frontier centroids, but in this work the candidate viewpoints are selected at centroids, with candidate views in four directions. This modified version of the multisensor frontier method is equivalent to the proposed approach if our method is not given any contextual information from the scan classifier.
@@ IV-B Simulation and Baseline Comparison
Name Difficulty [5] [41] Exploration [42] Classification (Our Method)
@@
Table I: Average time (s) for each method to find the object of interest in the given simulation environment.
@@
Two Gazebo world environments are tested with easy and hard setups in each. A figure depicting them is shown on the project website. The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES. The easy setting incorporates fewer objects in the environment while hard has more objects in the environment. The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position. The easy and hard settings are defined by the search target being in the same or different half of the environment as the robot starting position. Both difficulty levels in each map are repeated 10 times. Average detection times and success rates are reported in Table I. The time limit given to planners is 2 and 3 minutes in the apartment easy and hard settings then 3 and 5 minutes in the office easy and hard settings. Trials where the planner fails or the time limit is reached are reported as a failure and failed trials are reported as the maximum allotted time.
@@
Results indicate that LIVES, the Ground-Truth variant and the Multisensor Frontier Exploration method significantly outperform the RRT and NBVP baselines in all settings except for the hard office setting, where the RRT method performs similarly to LIVES and Ground-Truth methods. In the easy apartment setting we observe similar performance among the three planners. However in the hard setting, where LIVES and Ground-Truth perform nearly identically, there is roughly a 15% improvement with the addition of the knowledge of non-map points compared with the Multisensor method. There exists a large gap in the easy office setting with a near 30% improvement with the addition of contextual LiDAR information over the Multisensor variant. In the hard office setting, LIVES and Ground-Truth variant outperform the RRT method by 5-12%. Overall, the results indicate that the proposed method to inform the NBV planner with non-map information outperforms the baselines across the full spectrum of environments. We hypothesize this is due to the requirement for the exploration planners to exhaustively explore the environment, whereas the proposed method focuses on non-map points, allowing the agent to move on when the target is unlikely to be nearby. These results support hypotheses (1) and (2) that contextual LiDAR information significantly improves performance in the search task.
Figure 4: Ground truth maps of the two experimental environments. (a) shows the 15mx38m hallway environment and (b) shows the 20mx30m apartment.
Figure 5: A time lapse of a simple experiment scenario. In (a) the robot pose is a centered on the visual costmap. The global search map, frontier points and non-map points are also shown. The target object is behind the robot starting position. The result of the first planning step is also shown. The planner takes advantage of the concentration of non-map points behind the robot. Figure (b) demonstrates the moment of detection and localization of the target object (suitcase).
@@ IV-C Hardware Experiments
Experiments take place in a 20mx30m indoor apartment setting and a 15mx38m hallway setting, shown in Fig. 4. The environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy. A Boston Dynamics Spot robot with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera is utilized. The search target of interest is one or two small rolling suitcases and YOLOv5 [18] is used for detection of the object. Objects are localized by first converting PointCloud data into the optical frame and then by conversion to pixel coordinates using camera intrinsics. A time lapse of one experiment in the apartment is shown in Fig. 5 where the search target is situated behind the robot starting position. This example demonstrates the benefit of non-map information. In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the one on the left hand side. However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby. Finally, we compare the accuracy of the trained classifier in these out of distribution environments to ground truth information. The evaluated accuracy in the two out of distribution experimental environments are $84.0168\%$ in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment. These results demonstrate the robustness of the model in unseen environments, suggesting the model successfully identifies features common to many indoor environments. The success of LIVES onboard a real robot verify hypothesis (3). Videos of experiments, code and implementation details can be found on the project website https://sites.google.com/view/lives-2024/home.
@@ V Conclusions
We present a method for search in unseen environments using contextual information to inform planning. Results show LIVES to outperform several baseline methods in varied environments by 10-30%. Ablative studies on the map-free scan classifier are performed to validate the choice in architecture and training scheme. The proposed approach is deployed in an unseen environment to certify the learning method and visual planner for real-world scenarios. Ongoing work aims to extend the LIVES planner to multiple robots.

### v8

@@ III-E2 Viewpoint Selection
Candidate viewpoints are selected based on a computationally efficient heuristic utility function. High value viewpoints are those that are near the robot, allow inspection of many unknown map points and result in inspection of non-map points. This process is shown in Fig. 2(f). The utility functions is a weighted sum of the following: (1) Penalize distance from robot position to viewpoint, (2) Reward expected number of unknown cells discovered, (3) Reward viewpoints with frontiers near the path to the candidate viewpoint, and (4) Reward viewpoints that result in inspection of a scaled number non-map points. Scaling parameters on each term are tuned during experiments. Weight are set such that the agent is inclined to prioritize non-map points and those viewpoints that result in high number of expected cells discovered. The next waypoint is selected by sending the highest utility viewpoint, $x^{*}_{\text{next}}$, to the robot.
@@ IV Results
We test three hypotheses through our evaluation: (1) The proposed approach to inform an exploring agent with non-map features is effective at finding search targets, (2) LIVES reduces detection time compared against existing exploration algorithms, and (3) The proposed method can be deployed successfully on a real robot in unseen indoor environments.
@@
Name Difficulty [bircher2016receding] [rrt] Exploration [vidal2020multisensor] Classification (Our Method)
@@
Table I: Average time (s) for each method to find the object of interest in the given simulation environments.
IV-A LiDAR Scan Classification
We compare the learned classification policy to the ground truth classifier. The output of the policy is a vector of size $n_{i}$ normalized between -1 and 1 for each scan point. If a value is non-negative, the point belongs to $\mathcal{E}^{\text{map}}$. Otherwise the point belongs to $\mathcal{E}^{\text{non-map}}$. We express performance in terms of per-scan accuracy of the classified LiDAR scans from the policy vs. ground truth from map information. This metric for classified scan $s^{\text{class}}_{i}$ at time $i$ compared to ground truth $s^{\text{class, true}}_{i}$ is given by
IV-B Simulation and Baseline Comparison
@@
Simulations are performed in Gazebo with the Turtlebot [tbot]. Move base is used to navigate to the given viewpoints [navigation-ros]. Baselines are: 1) a modified Multisensor Frontier Exploration [vidal2020multisensor], 2) the Next-Best View Planner (NBVP) [bircher2016receding], 3) a Rapidly-Exploring Random Tree (RRT) [rrt], and 4) LIVES given ground-truth scan classifications. We make the following modifications to ref. [vidal2020multisensor] for our environments. The method is projected into 2D and viewpoints are sampled at centroids instead of surrounding them. This modified multisensor frontier method is equivalent to the proposed approach without segmented scan information. Two Gazebo worlds are tested with easy and hard setups in each. A figure depicting them is shown on the project website. The first map is the apartment (20x30m) and examines the impact of non-target objects on LIVES. The easy setting has fewer objects in the environment while hard has more objects. The second map is the office (25x45m) and tests the method in a large environment under significantly varied target position. Easy and hard settings are defined by the search object in the same or different half of the environment as the robot starting position. Both difficulty levels in each map are repeated 10 times. Average detection times and success rates are reported in Table I. The time limit given to planners is 2 and 3 minutes in the apartment easy and hard settings then 3 and 5 minutes in the office easy and hard settings. Trials where the planner fails or the time limit is reached are reported as a failure and failed trials are reported as the maximum allotted time.
@@
Results indicate that LIVES, the Ground-Truth variant and the Multisensor Frontier Exploration method significantly outperform the RRT and NBVP baselines in all settings except for the hard office setting, where the RRT method performs similarly to LIVES and Ground-Truth methods. In the easy apartment setting we observe similar performance among the three planners. In the hard setting LIVES and Ground-Truth perform nearly identically and show a 15% improvement with the addition of non-map points compared with the Multisensor method. There exists a large gap in the easy office setting with a near 30% improvement using contextual LiDAR information over the Multisensor variant. In the hard office setting, LIVES and Ground-Truth variant outperform the RRT method by 5-12%. The results indicate that LIVES outperforms the baselines across the full spectrum of environments. We hypothesize this is due to the requirement for the exploration planners to exhaustively explore the environment, whereas the proposed method focuses on non-map points, allowing the agent to move on when the target is unlikely to be nearby. These results support hypotheses (1) and (2) that contextual LiDAR information significantly improves performance in the search task.
@@ IV-C Hardware Experiments
Experiments are in a 20mx30m apartment and a 15mx38m hallway environment. These two environments are not included in dataset $\mathcal{D}$ used to train the pixel-wise classification policy. A Boston Dynamics Spot is equipped with a Velodyne Puck LiDAR and Azure Kinect RGB-D camera. The search object is one or two small rolling suitcases. YOLOv5 [yolov5] is used for detection. Objects are localized by first converting PointCloud data into the optical frame and then converting to pixel coordinates using camera intrinsics. Bounding boxes are then used to obtain the depth information from coinciding pixels. A time lapse of a simple trials in the apartment is shown in Fig. 5. The search target is situated behind the robot starting position. This example demonstrates the benefit of non-map information. In (b), if selecting simply based on number of unknown cells discovered, the agent may select another frontier, for example the candidate on the left hand side.
Figure 5: A time lapse of a simple scenario. The ground truth map (20x30m) is shown in the lower right corner (a). This is unavailable to the robot and not included in the training dataset. In (a), the global search map, frontier points and non-map points are also shown. The target object is behind the robot starting position. The result of the first planning step is also shown. The planner takes advantage of the concentration of non-map points behind the robot. Figure (b) demonstrates the moment of detection and localization of the target object (suitcase).
Figure 6: A time lapse of the hallway environment (15x38m) experiment scenario. In the lower right corner of (a) is the ground truth map with the suitcase locations and non-map points near the robot shown. This information is not in the training dataset and not given to the robot. In (a) the robot pose is a centered on the visual costmap. The first pose selected inspects the non-map points behind the robot. In (b) the robot chooses to turn left at the hallway split, recognizing a trash can as a non-map point. This correspondence can be seen between the non-map point in (b) appearing in the viewpoint visited by the robot in (c). In (d) the robot selects the circled viewpoint, facing to the left. This results in viewpoint (e), inspecting the suitcase, which was recognized as a non-map points in (d). In (f) the robot has just localized the first suitcase and takes the decision to return to the fork for non-map points found by that candidate viewpoint, leading to the second suitcase. (g) shows the overall robot path overlaid on the ground truth map.
However, equipped with this contextual information the agent selects the frontier that enables inspection of many non-map features nearby, resulting in detection. We compare the accuracy of the trained classifier in these out of distribution environments to ground truth information. The evaluated accuracy in the two out of distribution experimental environments are $84.0168\%$ in the apartment over a 105 second deployment and $84.869\%$ in the hallway over a 230 second deployment. These results demonstrate the robustness of the model in unseen environments, suggesting the model successfully identifies features common to many indoor environments. The success of LIVES onboard a real robot verify hypothesis (3). Videos of experiments, code and implementation details can be found on the project website https://sites.google.com/view/lives-2024/home. The video demonstrates three trials. The first and second take place in the apartment with one and two search objects, respectively. In both apartment scenarios, the robot tends to look towards furniture like couches, shelves and tables. Similarly, Spot also tends to inspect trash cans and other objects in the hallway. This behavior demonstrates the impact of non-map points and suggests the efficacy of the segmentation policy. Several key moments in the hallway search scenario are shown in the time lapse in Fig. 6.
@@ V Conclusions
We introduce LIVES, a method to leverage contextual information for search in unseen environments. It is shown to outperform baselines by 10-30% in simulation. Through ablative studies, we validate our architecture and training approach for map-free scan classification. Our experiments confirm the effectiveness of the visual search planner and learning method in real-world scenarios. Ongoing work plans to expand to multi-robot applications.
VI Acknowledgments
This research was supported in part by NSF Award #2219236 (GCR: Community Embedded Robotics: Understanding Sociotechnical Interactions with Long-term Autonomous Deployments) and Living and Working with Robots, a core research project of Good Systems, a UT Grand Challenge. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author and do not necessarily reflect the views of the National Science Foundation.
[…]